OpenStudy (anonymous):

Change to a simplified fraction or mixed number. 1 8 - % 5

OpenStudy (anonymous):

41/5 %

OpenStudy (anonymous):

it is only giving me 41/50 or 41/500

OpenStudy (anonymous):

i dont know what the correct one is

OpenStudy (anonymous):

and 41/100

OpenStudy (anonymous):

8 1/5 % = 8.2% = 0.082 % = 82/1000 = 41/500

OpenStudy (anonymous):

ok thank you